OpenStudy (anonymous):

Solve for y: 2/3y + 4 = 6 y = 3 y = 7 y = -3 y = -7

OpenStudy (anonymous):

@jim_thompson5910 can you help with this?

jimthompson5910 (jim_thompson5910):

2/3y + 4 = 6 2y + 12 = 18 ... multiply everything by 3 2y = 18 - 12 2y = 6 y = ???

OpenStudy (anonymous):

y=3?@jim_thompson5910

OpenStudy (anonymous):

Yes. y = 3

jimthompson5910 (jim_thompson5910):

yep, y = 3